# SAM-LAD: Segment Anything Model Meets Zero-Shot Logic Anomaly Detection

### v2

@@ IV-D4 DigitAnatomy
IV-E Further Analysis
The outstanding logical anomaly detection results of our framework are based on the good performance of the Object Matching Model (OMM). Therefore, to verify the reasons behind the exceptional performance, for each category in MVTec LOCO AD and DigitAnatomy dataset, we compute the precision, recall, and F1 score of OMM’s matching results, which are shown in Table V.
The results demonstrate the OMM’s strong matching capability and validate its effectiveness. However, the matching accuracy for the pushpins and splicing connectors categories drops significantly. This occurs because each object is very similar in these scenarios, posing a challenge for the OMM. Instead of using unmatched object masks as the anomaly score, we designed the AMM (Anomaly Measurement Module) to further detect anomalies. This design reduces the SAM-LAD’s dependency on matching accuracy, as shown by the final anomaly detection results in Section IV-D.
TABLE V: The Performance of the OMM’s Matching capability in MVTec LOCO AD dataset and Digitanatomy Dataset.
@@ IV-E Further Analysis
IV-F Ablation Experiment
IV-F1 Impact of the Backbone
We use the pre-trained backbone as the feature extractor in the proposed SAM-LAD. In this study, we primarily considered three different pre-trained backbone networks: the Wide-ResNet50 of the CNN type, the Swin-transformer of the ViT type, and the DINOv2 of the ViT type. The main results are presented in Table VI, where it is distinctly that DINOv2 manifests the most superior performance. Therefore, we adopt DINOv2 as the feature extractor for the proposed framework.
TABLE VI: Ablation Experiments of the Different Backbone on the MVTec LOCO AD Dataset.
@@ IV-F1 Impact of the Backbone
IV-F2 Impact of the FeatUp Configuration
@@
Figure 13: Visualizing PCA components with different upsampling configurations of the FeatUp.
The main results are shown in Table VII. Compared to the original feature maps, FeatUp significantly enhances detection capabilities. Specifically, as the upsampling factor increases, SAM-LAD’s detection and segmentation performance improves. However, upsampling the feature maps by 16$\times$ and 32$\times$ greatly increases the model’s FLOPs and drastically reduces inference speed, which is impractical for industrial use. Therefore, to balance detection performance with real-time inference needs, we chose an 8$\times$ upsampler to restore lost spatial information in the feature maps for subsequent computations.
TABLE VII: Ablation experiments of the Different FeatUp’s Upsampling Factors on the MVTec LOCO AD Dataset.
@@
Efficency FLOPs(Gb) 9.2 15.2 23.8 34.6 65.1 94.2
FPS 7.2 5.5 3.4 1.9 0.9 0.1
TABLE VIII: Ablation Experiments of the Different Feature Compress Methods on the MVTec LOCO AD Dataset.
@@ IV-F3 Impact of the DCGA
To validate the effectiveness of DCGA, we conducted an ablation experiment. DCGA was designed to extract high-dimensional features into a single vector, simplifying the similarity computation between objects. Pooling is the most widely used method for feature extraction. We compared DCGA with global average pooling (GAP) and global max pooling (GMP), as shown in Table VIII. GMP outperforms GAP because only certain areas of the object feature map have discernible values, while the background is mostly null. Therefore, GMP is more effective in capturing object features in such scenarios. However, simply selecting the maximum value for pooling is insufficient. The introduction of DCGA greatly enriched the description of object features. Using graph neural network principles, DCGA extracts effective object features from high-dimensional channels and captures feature interrelationships among objects in the same scene. This substantially improves the matching capabilities of the OMM, confirming the validity of DCGA.
V Conclusion
@@
In this paper, we novelty propose a zero-shot framework called SAM-LAD to address logical anomaly detection in complex scenes. We introduce a pre-trained SAM to obtain masks for all objects in the query image. Utilizing the pre-trained DINOv2 and FeatUp operations, we derive the upsampled feature map. Through nearest neighbor search on the query image, we obtain a reference image and its corresponding upsampled feature map. By sequentially combining object masks with upsampled feature maps, we acquire each object feature maps in both the query and reference images. Subsequently, we regard each object as a key point and employed the proposed DCGA mechanism to efficiently compress each object’s feature map into a feature vector. Then, we propose the OMM, matching all object feature vectors from the query image with those in the reference image to obtain a matching matrix. Finally, based on the matching matrix, we propose AMM to compute the distribution difference estimation of the matched objects’ features, resulting in the final anomaly score map.
Nevertheless, within the scope of our investigation, a limitation has come to light, that could potentially impede the overall performance of our framework. Specifically, when objects in the scene all belong to the same category, due to their high similarity, the object matching model exhibit some false positives, such as the pushpins shown in Table V. Consequently, as part of our future work, we aim to explore potential enhancements to mitigate these issues and further boost the performance of our framework.

### v4

@@ 4.4.4 DigitAnatomy
4.5 Further Analysis
The outstanding logical anomaly detection results of our framework are based on the excellent performance of the Object Matching Model (OMM). Therefore, to verify the reasons behind the exceptional performance, for each category in MVTec LOCO AD and DigitAnatomy dataset, we compute the precision, recall, and F1 score of OMM’s matching results, which are shown in Table 5.
The results demonstrate the OMM’s strong matching capability and validate its effectiveness. However, the matching accuracy for the pushpins and splicing connectors categories drops significantly. This occurs because each object is very similar in these scenarios, posing a challenge for the OMM. Instead of using unmatched object masks as the anomaly score, we designed the AMM (Anomaly Measurement Module) to further detect anomalies. This design reduces the SAM-LAD’s dependency on matching accuracy, as shown by the final anomaly detection results in Section 4.4.
Table 5: The Performance of the OMM’s Matching capability in MVTec LOCO AD dataset and Digitanatomy Dataset.
@@ 4.5 Further Analysis
4.6 Ablation Experiment
4.6.1 Impact of the Backbone
We use the pre-trained backbone as the feature extractor in the proposed SAM-LAD. In this study, we primarily considered three different pre-trained backbone networks: the Wide-ResNet50 of the CNN type, the Swin-transformer of the ViT type, and the DINOv2 of the ViT type. The main results are presented in Table 6, where it is clear that DINOv2 manifests the most superior performance. Therefore, we adopt DINOv2 as the feature extractor for the proposed framework.
Table 6: Ablation Experiments of the Different Backbone on the MVTec LOCO AD Dataset.
@@ 4.6.1 Impact of the Backbone
4.6.2 Impact of the FeatUp Configuration
@@
Figure 13: Visualizing PCA components with different upsampling configurations of the FeatUp in MVTec LOCO AD.
The main results are shown in Table 7. Compared to the original feature maps, FeatUp significantly enhances detection capabilities. Specifically, as the upsampling factor increases, SAM-LAD’s detection and segmentation performance improves. However, upsampling the feature maps by 16$\times$ and 32$\times$ significantly increases the model’s FLOPs and drastically reduces inference speed, which is impractical for industrial use. Therefore, to balance detection performance with real-time inference needs, we chose an 8$\times$ upsampler to restore lost spatial information in the feature maps for subsequent computations.
Table 7: Ablation experiments of the Different FeatUp’s Upsampling Factors on the MVTec LOCO AD Dataset.
@@
Efficency FLOPs(Gb) 9.2 15.2 23.8 54.7 165.1 294.2
FPS 87.2 45.5 21.4 8.9 2.9 1.1
Table 8: Ablation Experiments of the Different Feature Compress Methods on the MVTec LOCO AD Dataset.
@@ 4.6.3 Impact of the DCGA
To validate the effectiveness of DCGA, we conducted an ablation experiment. DCGA was designed to extract high-dimensional features into a single vector, simplifying the similarity computation between objects. Pooling is the most widely used method for feature extraction. We compared DCGA with global average pooling (GAP) and global max pooling (GMP), as shown in Table 8. GMP outperforms GAP because only certain areas of the object feature map have discernible values, while the background is primarily null. Therefore, GMP is more effective in capturing object features in such scenarios. However, simply selecting the maximum value for pooling is insufficient. The introduction of DCGA greatly enriched the description of object features. Using graph neural network principles, DCGA extracts compelling object features from high-dimensional channels and captures feature interrelationships among objects in the same scene. This substantially improves the matching capabilities of the OMM, confirming the validity of DCGA.
5 Discussion
Within the scope of our investigation, we have identified a limitation that may hinder the overall performance of SAM-LAD. Specifically, the SAM-LAD encounters limitations when multiple objects of the same category are present in the scene to be inspected. This limitation stems from the fact that the core of SAM-LAD relies on an explicit matching principle to identify unmatched objects and detect anomalies. For instance, in the case of the logical anomaly in the box of pushpins shown in Fig.14(a), where each compartment is expected to contain only one pushpin, SAM-LAD is expected to detect the abnormality of having two pushpins in each compartment. Unfortunately, due to the high similarity of all pushpin features in the scene, the OMM of SAM-LAD fails to accurately identify the extra pushpins in each compartment when compared with the normal reference image. As a result, as shown in Fig.14(b), SAM-LAD fails to detect the logical anomaly in this scenario. The core issue lies in SAM’s difficulty with segmentation granularity, often resulting in outputs that are either overly fine or too coarse.
Figure 14: Failure case of SAM-LAD in MVTec LOCO AD.
Future research could integrate state-of-the-art visual foundation models and clustering techniques to achieve precise segmentation of key contextual objects in a scene. Specifically, the Recognize Anything Model[53] could be employed to identify objects in the scene and generate corresponding labels. Subsequently, the Grounded SAM[54] method generates masks for all detected elements. If Grounded SAM produces multiple masks, indicating the presence of multiple objects, clustering methods will be applied to refine these masks. Finally, the refined segmentation mask of each key object is passed to OMM. Even for objects of the same type, OMM can accurately match them due to the refined segmentation map.
Another future promising approach is to fine-tune SAM within a specific domain to enhance its segmentation performance in targeted application scenarios, such as industrial inspection. Although SAM is a segmentation model with robust generalization capabilities, its performance can still be further optimized for specific domains through fine-tuning. Specifically, industrial components often exhibit repetitive patterns, distinct textures, and unique visual features that differ significantly from the general datasets used to train SAM. Fine-tuning domain-specific data can enable SAM to adapt to these features, thereby achieving more accurate and consistent segmentation results. This directly addresses the challenge where SAM-LAD struggles to accurately segment highly similar objects in a scene, which weakens anomaly detection performance. Additionally, industrial inspection often involves controlled yet varying lighting conditions. A fine-tuned SAM model can adapt to these conditions and mitigate the impact of lighting variations on segmentation performance.
6 Conclusion
@@
In this paper, we novelty propose a zero-shot framework called SAM-LAD to address logical anomaly detection in complex scenes. We introduce a pre-trained SAM to obtain masks for all objects in the query image. Utilizing the pre-trained DINOv2 and FeatUp operations, we derive the upsampled feature map. Through the imageNNS on the query image, we obtain a reference image and its corresponding upsampled feature map. By sequentially combining object masks with upsampled feature maps, we acquire each object feature map in both the query and reference images. Subsequently, we regarded each object as a key point and employed the proposed DCGA mechanism to efficiently compress each object’s feature map into a feature vector. Then, we propose the OMM, matching all object feature vectors from the query image with those in the reference image to obtain a matching matrix. Finally, based on the matching matrix, we propose AMM to compute the distribution difference estimation of the matched objects’ features, resulting in the final anomaly score map.